OpenStudy (anonymous):

discuss the convergence of Xn = root 6Xn where X1=1

OpenStudy (anonymous):

\[x_{n} =\sqrt{6x_n} where x_1=1\]

OpenStudy (anonymous):

@dan815 any clue?

OpenStudy (dan815):

no sorry i dont even get the question

OpenStudy (dan815):

@ganeshie8

OpenStudy (anonymous):

oh ok,noprobs

OpenStudy (anonymous):

@BSwan

OpenStudy (anonymous):

i think there is a typo, xn=sqrt 6xn does not seem to be right x1=sqrt 6 x1=sqrt 6 \(\neq\) plz rewrite the formula and make sure to type the correct one :)

OpenStudy (anonymous):

it is correct

OpenStudy (anonymous):

well then sorry i cant help :)

OpenStudy (anonymous):

well,no probs

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@amistre64 @Hero

OpenStudy (anonymous):

I have to agree with @BSwan, if this is supposed to be a recursive sequence, it should say something like \[x_{n-1}=\sqrt{6x_n}\] or \[x_{n}=\sqrt{6x_{n-1}}\]

OpenStudy (anonymous):

:( its not like that in question paper

OpenStudy (amistre64):

can you send a link/screenshot/pic of the question?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

yeah, thats on odd question. the pictures too fuzzy but it looks like what youve posted in general so .... just going off of whats there (xn) = sqrt(6xn) (xn)^2 - 6(xn) = 0 if x1=1 we have a false statement still doesnt make alot of sense yet. any examples in your material?

OpenStudy (anonymous):

i was thinking of something like this xn=6 xD so its constant that would be converge , but x1 make a contradiction to me lol

OpenStudy (anonymous):

i only hv questions -.- i was absent during this whole class,so am having to study it here as the teachers say they wont be teaching it again for me

OpenStudy (anonymous):

wow ! and u cant ask them anything ??

OpenStudy (anonymous):

appearntly i cant,they insult us in front of other teachers if i go n ask,its not realy a good idea to ask these A-holes :P

OpenStudy (anonymous):

haha lolz xD will i think ur smart :)

OpenStudy (anonymous):

you think? am honored? :D

OpenStudy (anonymous):

yep :)